Tohban Report 2013-09-11

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 04 Sep 2013
End Date: 11 Sep 2013
Tohban: Nicole Duncan
Tohban email:
Next Tohban: Hazel Bain
List all reports


Solar Activity

Solar activity was very low this week, there haven't been any flares in the past two days. There are currently two active regions on disk, neither of which have produced any flares to date. Stereo B shows at least one active region to rotate on disk in the next week.

How many GOES flares occurred?

 Flares above B, C, M, X class were     17    12     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      9     6     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      6     3     0     0

There were RHESSI flares/GOES flares 173 / 29 over the time range 03-Sep-13 10-Sep-13

Memory Management

The SSR is emptying at the end of the pass set. Maximum reached this week was 22%.

Spacecraft Status

Spacecraft status is nominal. The cold plate 1 and 2 are at 116.8 K and 115.2 K respectively. ~1 deg decrease from last week, as expected from changing eclipse length. Cryocooler nominal 13mG. Spin 14.8rpm.

Data Gaps

All filled in.

Day 248, 20second issue due to issue with ApID 102.

Detector issues

No additional D2 or D1 spikes were seen.

There were 3 spikes that occured throughout the detectors and also appeared in the particle rates, suggesting they were the result of a corrupted packet (not an actual spacecraft issue). These occurred on Sept 5th at 05:36:10UT and on Sept 7th at 00:53:58UT and 1:11:30UT.

The D9 rear segment noise returned this week. The brief dropouts grew in strength, reaching a maximum on the 9th. Their severity was no greater than in the past and there were no tangible consequences to SSR or livetime, so no action was taken. The severity and duration has begun to decrease and as of the 10th, was comparable to those seen on the 4th.

Other notes

A thermistor on PRT 3, solar array 2 is failing. When it permanently fails, the result will be that it just reads high. It hits the rail at 127C, during every orbit. Its gone through 67,000 temperature cycles in its lifetime.

Spacecraft Management

No spin-ups or special commanding this week.

Decimation Normal/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Enabled, during entire eclipse
Night time data (rears) Enabled, during entire eclipse
Require extra passes? No?
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Increased noise in Detector 7 continues.
Personal tools